Circular No. 7695

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q4 (NEAT) S. H. Pravdo, E. F. Helin, and K. J. Lawrence, Jet Propulsion Labora-

tory, report the discovery of a comet on CCD images taken with the 1.2-m Schmidt telescope at Palomar in the course of the NEAT program; their images on Aug. 24, 26, and 27 show a spherically symmetrical nebulosity with diameter ∼ 8′′. CCD total magnitude estimates: Aug. 27.08 UT, 17.8 (J. Tichá, M. Tichý, and P. Jelinek, Kleť Observatory, 0.57-m reflector); Aug. 27.47, 17.3 (P. J. Shelus, McDonald Observatory, 0.76-m reflector).

2001 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 24.40491  2^{h}44^{m}36.^{s}94  −24^{o}58′42′′.0  20.0

Full astrometry is given on MPEC 2001-Q67. The orbital elements are very uncertain, and those by B. G. Marsden given below represent a compromise solution:

T = 2005 Aug. 25.340 TT  ω = 325.^{o}884
Ω = 218.402 }2000.0
q = 4.09424 AU  i = 108.688
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 19 2^{h}45.^{m}08 −24^{o}31.^{′}8 11.099 11.495 110.^{o}6 4.^{o}7 17.3
29 2 44.07 −25 21.9 10.930 11.437 117.9 4.5 17.3
Sept. 8 2 42.42 −26 13.0 10.777 11.380 124.6 4.2 17.2
18 2 40.12 −27 03.9 10.643 11.322 130.6 3.9 17.2
28 2 37.24 −27 52.9 10.531 11.264 135.2 3.6 17.1

V4739 SAGITTARII = NOVA SAGITTARII 2001 No. 2 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4739 Sgr has been given to the nova reported on IAUC 7692.

K. Takamizawa, Saku-machi, Nagano, Japan, reports prediscovery limiting magnitudes (nova not detected) from unfiltered T-Max 400 films: Aug. 24.477 UT, [12.1; 24.480, [12.1.

A. C. Gilmore forwards additional photometry by P. M. Kilmartin at Mt. John: Aug. 27.408 UT, V = 7.66, U − B = −0.78, B − V = +0.55, V − R = +0.60, V − I = +1.10, air mass 1.06; 27.506, 7.94, −0.86, +0.53, +0.68, +1.22, 1.35; 27.552, 8.04, −0.87, +0.53, +0.72, +1.27, 1.71; 27.615, 8.16, −0.90, +0.53, +0.77, +1.33, 2.89; 28.318, 9.12, −0.93, +0.28, +1.09, +1.53, 1.06; 28.366, 9.27, −0.86, +0.26, +1.13, +1.56, 1.03.

2001 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green